# bHUN-REN Wigner Research Centre for Physics, Konkoly-Thege Miklos ut 29-33, 1121 Budapest, Hungary

### v1

@@ 1 Introduction and summary
    For $\alpha=0$, only the off-diagonal blocks in (4) vanish, and $Z(g)$ can be written as a product of the determinants of the diagonal blocks. For different forms of the symbol, and values of $\ell$, these determinants describe several observables in planar $\mathcal{N}=4$ SYM, for instance, the flux tube correlators Beisert et al. (2007); Belitsky and Korchemsky (2020a); Basso et al. (2020) and correlation functions of infinitely heavy half-BPS operators Coronado (2019, 2020); Kostov et al. (2019a); Bargheer et al. (2019a); Kostov et al. (2019b); Belitsky and Korchemsky (2020a); Bargheer et al. (2019b). For other choices of the symbol, they also give the leading non-planar correction to the partition function of $\mathcal{N}=2$ SYM Beccaria et al. (2020, 2021b, 2021a, 2022, 2023), and in another case coincide with the Tracy-Widom distribution that describes the eigenvalue distribution in the Laguerre ensemble near the hard edge Tracy and Widom (1994).
@@
As discussed in Bajnok et al. (2025b), the determinant (1) can be reformulated in the form of a Fredholm determinant of an integral operator whose kernel can be written as a two-by-two matrix of Bessel kernels (for this reason, I simply refer to (1) as a determinant with a matrix Bessel kernel). In recent years a huge development was made in the study of Fredholm determinants with Bessel kernel. Based on previous results from the mathematical literature (see, for instance, (Szegő, 1915, 1952; Kac, 1954; Akhiezer, 1966; Fisher and Hartwig, 1969; Böttcher, 1995; Basor, 2012; Its et al., 1990b; Korepin et al., 1993; Tracy and Widom, 1994)), it was shown that these determinants and the resolvents of the Bessel kernel satisfy a system of coupled integro-differential equations and they provide a systematic way to expand them both for weak and strong coupling in an efficient way (Its et al., 1990a; Belitsky and Korchemsky, 2020b, 2021; Bajnok et al., 2024, 2025c, 2025b). By high order numerical computations, it was shown that these series are asymptotic and a resurgence analysis (Mariño, 2014; Dorigoni, 2019; Aniceto et al., 2019) is essential to obtain a finite physical answer. The resurgence properties for the weak and strong coupling regimes were studied in Basso and Korchemsky (2009); Aniceto (2016); Dorigoni and Hatsuda (2015); Bajnok et al. (2024, 2025c, 2025b); Dunne (2025); Bajnok et al. (2025d). Relations for the determinant (1) for different values of the parameter $\ell$ were also found in Korchemsky (2025).
In this paper, I continue the program of Bajnok et al. (2024, 2025c, 2025b, 2025d) and investigate further properties of the strong-coupling expansion of (1) for arbitrary values of $\alpha$ and $\ell$, with the specific symbol (7). In Bajnok et al. (2025b) we have already presented a method to compute the large-$g$ expansion for $Z_{\ell}(g)$ in this case up to arbitrary orders in the transseries parameters and investigated its resurgence properties. Our method was based on first computing the logarithm of the determinant $\mathcal{F}_{\ell}(g)$ through a set of integro-differential equations and then exponentiating the resulting series to obtain the strong coupling expansion of $Z_{\ell}(g)=e^{\mathcal{F}_{\ell}(g)}$.
@@
In our latest paper Bajnok et al. (2025d), we have shown that in case of $\alpha=0$ (which we studied previously in Bajnok et al. (2024) and Bajnok et al. (2025c)), there is a simple underlying structure in the large $g$ expansion of the determinant itself. We found that the non-perturbative contributions appear only at first orders in certain exponential scales, which are related to the zeros of the function $1-\chi(x)$. Furthermore, these contributions can be easily obtained from the perturbative sector by applying simple transformations on the coefficients to the perturbative $1/g$ series. We also found that in this new structure the resurgence between different non-perturbative sectors appears in a natural way.
The question then arises whether the same principle can be generalized to observables given by the matrix Bessel kernel defined in (4) with symbol (7) and arbitrary values of $\alpha$. In this paper, I demonstrate that, by rearranging the strong coupling expansion found in Bajnok et al. (2025b), the results of Bajnok et al. (2025d) can be extended to the observables discussed in points 2. and 3. above, making it extremely efficient to generate the complete transseries for these physical quantities as well. This new form of the transseries also provides the opportunity to give a complete description to the resurgence structure of the strong coupling expansion of these observables.
@@
to make the dependence on the moments $\mathcal{I}_{n}$ explicit. With the system of integro-differential equations studied in Bajnok et al. (2025b), the $1/g$ expansion of $\mathcal{D}\left[\mathcal{I}_{n}\right](g)$ can be efficiently generated both analytically and numerically. Then applying the same ideas as in Bajnok et al. (2025d), I show that the exponentially small corrections can be easily generated from the perturbative functions with the simple rule:
@@
The advantage of writing the transseries as (52) instead of (9), is that the exponential scales are in one-to-one correspondence with the zeros of the relevant symbol (11). Moreover, the exponentially small corrections are at most first order in every scale $e^{-8\pi gx_{l}^{+}}$ and $e^{-8\pi gx_{j}^{-}}$, and the corresponding $1/g$ expansions can be easily obtained from the perturbative sector with the simple rule in (16). This is the same structure as was studied in Bajnok et al. (2025d). Therefore, (52) puts the observables in points 2. and 3. defined through the matrix Bessel kernel with different values of the mixing angle $\alpha$ in the universal strong coupling structure of Bajnok et al. (2025d). It also provides a more natural framework for describing the resurgence properties of the above-mentioned observables at large values of the ’t Hooft coupling.
According to the results of this paper together with Bajnok et al. (2025d), there is a simple underlying structure in the strong coupling expansion for certain observables of $\mathcal{N}=4$ SYM, at least for those given by the determinant (1). It is an open question whether the same structure holds for other observables in $\mathcal{N}=4$ SYM as well, or physical quantities in different but related models, such as the $O(6)$ model Polyakov (1980); Korchemsky and Radyushkin (1987); Alday and Maldacena (2007). It would also be interesting to relate the obtained transseries structure to the results of Bajnok et al. (2023, 2025a) where a complete resurgence analysis was carried out for a family of generalized energy densities in the $O(N)$ models.
@@ 2 Strong coupling expansion
$\mathcal{I}_{n}=(-1)^{n}\mathcal{I}_{n}|_{a\to-a}$ (21)
@@
$\mathcal{Z}^{(0,1)}(g)=\left.\mathcal{Z}^{(1,0)}(g)\right|_{a\to-a}\,$, (28)
and the same property for the corresponding exponential scales $\Lambda^{2}_{-}$ and $\Lambda^{2}_{+}$. Therefore the prefactor $P^{(0,1)}$ and the coefficients $f^{(0,1)}_{k}$ can be easily obtained from $P^{(1,0)}$ and $f^{(1,0)}_{k}$ by changing $a\to-a$.
@@
$\mathcal{Z}^{(n,m)}(g)=\left.\mathcal{Z}^{(m,n)}(g)\right|_{a\to-a}\,.$ (34)
The exponential scales in (10) also respect this symmetry.
@@
$\mathcal{I}_{n}=(-1)^{n}\left.\mathcal{I}_{n}\right|_{-a}=\left.\mathcal{I}_{n}\right|_{a+1}=(-1)^{n}\left.\mathcal{I}_{n}\right|_{1-a}\,.$ (40)
@@
$\left.\mathcal{I}_{n}\right|_{a\to a+m}\equiv\mathcal{I}_{n}\,$, (41)
@@
That is, the large-$g$ expansion is the same as in the principal branch with the same moments given in (20), but the explicit dependence on $a$ is shifted according to the branch of $a$. Later we will see, that the same property appears in the relation connecting the perturbative series with the non-perturbative corrections.
One final remark is that although it is formally possible to determine the transseries (9) up to arbitrary order in $1/g$ and $\Lambda_{\pm}^{2}$ with our previous method established in Bajnok et al. (2025b), however, due to the scaling (36), beyond a certain exponential order, it becomes difficult to gain analytical results for the strong coupling coefficients. Following the idea of Bajnok et al. (2025d), in the next section I will show, that (9) can be recast into a new form, in which a natural connection between the perturbative and non-perturbative contributions emerges, making the computation of non-perturbative corrections extremely efficient.
@@ 3 The transseries structure from a different angle
The reason for the cancelation of terms that do not satisfy the condition in (35) remained unknown, as it does not follow trivially from the method presented in Bajnok et al. (2025b). It can also be seen that, by applying certain transformations to the parameter $a$ and the integrals $\mathcal{I}_{n}$, the non-perturbative functions $\mathcal{Z}^{(n,m)}(g)$ with $(n,m)\neq(0,0)$ can easily be obtained from the perturbative part $\mathcal{Z}^{(0,0)}(g)$. In the following I will show that the procedure of Bajnok et al. (2025d) can be generalized to the matrix Bessel kernel (4) with symbol (7), and these properties follow automatically from the analytic structure of the modified symbol (11) and its successive redefinition at each exponential level.
@@
From the method of Bajnok et al. (2025b) and the structure studied in Bajnok et al. (2025d), we expect the appearance of non-perturbative contributions at orders of $e^{-8\pi gx^{\pm}_{j}}$. If we look at the exponential scales $\Lambda_{-}^{2n}\Lambda_{+}^{2m}$ multiplying $\mathcal{Z}^{(n,m)}(g)$ in the transseries (9), we find that the exponential prefactors are:
@@
For different values of $n$ and $m$, the exponent contains both +a+a and a-a terms and can be written as finite linear combinations of the zeros $x_{j}^{\pm}$, with each $x_{j}^{\pm}$ appearing with a multiplicity of one. Therefore, together with the findings of Bajnok et al. (2025d), this suggests that in the case of matrix Bessel kernel, non-perturbative contributions contain exponential factors that are at most first order in each parameter $e^{-8\pi gx_{j}^{+}}$ and $e^{-8\pi gx_{j}^{-}}$. Therefore, the exponential orders can be parametrized as:
@@
$S^{(\delta^{+},\delta^{-})}$ $=S^{(\delta^{-},\delta^{+})}\left.\right|_{a\to-a}\,$, (54)
$\mathcal{D}^{(\delta^{+},\delta^{-})}(g)$ $=\mathcal{D}^{(\delta^{-},\delta^{+})}(g)\left.\right|_{a\to-a}\,.$ (55)
@@ 3.1 $1/g$ expansion
Now, this function has zeros at $x=2i\pi x_{j}^{-}$ with $j\geq 1$ and poles at $x=2i\pi x^{-}_{0}$ and $x=2i\pi y_{j}$ with $j\geq 0$. These transformation rules are analogous to what we have found in Bajnok et al. (2025d).
@@ 3.3 A practical example
with $\zeta(x)$ is the Riemann zeta function and $\beta(x)$ is the Dirichlet beta function. The same special functions already appeared in the analysis of the strong coupling expansion of the cusp anomalous dimension. For example, in equations (28) and (29) of Basso et al. (2008) the perturbative part of the cusp anomalous dimension were given in terms of some coefficients $c_{n}$ proportional to $\zeta(x)$ and $\beta(x)$ with odd and even integer arguments respectively. Hence these coefficients are in one-to-one correspondence with moments $\mathcal{I}_{n}$, namely:
@@
The perturbative part in (3.3) reproduces the expansion presented in Basso et al. (2008), while the first non-perturbative correction coincides with the results of Basso and Korchemsky (2009). With the method discussed above, any higher order non-perturbative contribution to the cusp anomalous dimension can be easily obtained in an extremely efficient way.
@@ 4.1 Stokes automorphism
$\mathcal{S}_{\pm}\left[\mathcal{D}^{(\{\},\{\})}\right](\tilde{g})=\tilde{g}\int_{0}^{\infty e^{i\epsilon}}dse^{-\tilde{g}s}\mathcal{B}\left[\mathcal{D}^{(\{\},\{\})}\right](s)\,.$ (108)
@@ 4.2 Asymptotic analysis
Notice that the leading asymptotic behavior of $d^{(\{\},\{\})}$, $d^{(\{0\},\{\})}$ and $d^{(\{\},\{0\})}$ with the coefficients given in (4.2) is compatible with results found in Bajnok et al. (2025b) for the functions $\mathcal{Z}^{(0,0)}(g)$, $\mathcal{Z}^{(1,0)}(g)$ and $\mathcal{Z}^{(0,1)}(g)$, however there are sign differences in the coefficients $A_{0}^{+}$ and $A_{0}^{-}$. As a verification of (4.2), I also performed a Richardson extrapolation on the coefficients of the perturbative function. Its asymptotic behavior agreed with (4.2) with the same coefficient $A_{0}^{+}$ as in (4.2). Furthermore, with the method discussed in Appendix G of Bajnok et al. (2025d), I was also able to separate the first subleading cut of $\mathcal{B}\left[\mathcal{D}^{(\{\},\{\})}\right](s)$. I found that around this cut, its asymptotic behavior is given by:

### v2

@@ 1 Introduction and summary
    For $\alpha=0$, the off-diagonal blocks in (4) vanish, and $Z(g)$ can be written as a product of the determinants of the diagonal blocks. For different forms of the symbol, and values of $\ell$, these determinants describe several observables in planar $\mathcal{N}=4$ SYM, for instance, the flux tube correlators Beisert et al. (2007); Belitsky and Korchemsky (2020a); Basso et al. (2020) and correlation functions of infinitely heavy half-BPS operators Coronado (2019, 2020); Kostov et al. (2019a); Bargheer et al. (2019a); Kostov et al. (2019b); Belitsky and Korchemsky (2020a); Bargheer et al. (2019b). For other choices of the symbol, they also give the leading non-planar correction to the partition function of $\mathcal{N}=2$ SYM Beccaria et al. (2020, 2021b, 2021a, 2022, 2023) and appear in the study of correlation functions in $\mathcal{N}=2$ superconformal quiver gauge theories Billo et al. (2021); Korchemsky and Testa (2025). In another case, they coincide with the Tracy-Widom distribution that describes the eigenvalue distribution in the Laguerre ensemble near the hard edge Tracy and Widom (1994).
@@
As discussed in Bajnok et al. (2025b), the determinant (1) can be reformulated in the form of a Fredholm determinant of an integral operator whose kernel can be written as a two-by-two matrix of Bessel kernels (for this reason, I simply refer to (1) as a determinant with a matrix Bessel kernel). In recent years a huge development was made in the study of Fredholm determinants with Bessel kernel. Based on previous results from the mathematical literature (see, for instance, (Szegő, 1915, 1952; Kac, 1954; Akhiezer, 1966; Fisher and Hartwig, 1969; Böttcher, 1995; Basor, 2012; Its et al., 1990b; Korepin et al., 1993; Tracy and Widom, 1994)), it was shown that these determinants and the resolvents of the Bessel kernel satisfy a system of coupled integro-differential equations and they provide a systematic way to expand them both for weak and strong coupling in an efficient way (Its et al., 1990a; Belitsky and Korchemsky, 2020b, 2021; Bajnok et al., 2024, 2025c, 2025b). By high order numerical computations, it was shown that these series are asymptotic and a resurgence analysis (Mariño, 2014; Dorigoni, 2019; Aniceto et al., 2019) is essential to obtain a finite physical answer. The resurgence properties for the weak and strong coupling regimes were studied in Basso and Korchemsky (2009); Aniceto (2016); Dorigoni and Hatsuda (2015); Bajnok et al. (2024, 2025c, 2025b); Dunne (2025); Bajnok et al. (2026b). Relations for the determinant (1) for different values of the parameter $\ell$ were also found in Korchemsky (2025).
In this paper, I continue the program of Bajnok et al. (2024, 2025c, 2025b, 2026b) and investigate further properties of the strong-coupling expansion of (1) for arbitrary values of $\alpha$ and $\ell$, with the specific symbol (7). In Bajnok et al. (2025b) we have already presented a method to compute the large-$g$ expansion for $Z_{\ell}(g)$ in this case up to arbitrary orders in the transseries parameters and investigated its resurgence properties. Our method was based on first computing the logarithm of the determinant $\mathcal{F}_{\ell}(g)$ through a set of integro-differential equations and then exponentiating the resulting series to obtain the strong coupling expansion of $Z_{\ell}(g)=e^{\mathcal{F}_{\ell}(g)}$.
@@
In our latest paper Bajnok et al. (2026b), we have shown that in case of $\alpha=0$ (which we studied previously in Bajnok et al. (2024) and Bajnok et al. (2025c)), there is a simple underlying structure in the large $g$ expansion of the determinant itself. We found that the non-perturbative contributions appear only at first orders in certain exponential scales, which are related to the zeros of the function $1-\chi(x)$. Furthermore, these contributions can be easily obtained from the perturbative sector by applying simple transformations on the coefficients to the perturbative $1/g$ series. We also found that in this new structure the resurgence between different non-perturbative sectors appears in a natural way.
The question then arises whether the same principle can be generalized to observables given by the matrix Bessel kernel defined in (4) with symbol (7) and arbitrary values of $\alpha$. In this paper, I demonstrate that, by rearranging the strong coupling expansion found in Bajnok et al. (2025b), the results of Bajnok et al. (2026b) can be extended to the observables discussed in points 2. and 3. above, making it extremely efficient to generate the complete transseries for these physical quantities as well. This new form of the transseries also provides the opportunity to give a complete description to the resurgence structure of the strong coupling expansion of these observables.
@@
to make the dependence on the moments $\mathcal{I}_{n}$ explicit. With the system of integro-differential equations studied in Bajnok et al. (2025b), the $1/g$ expansion of $\mathcal{D}\left[\mathcal{I}_{n}\right](g)$ can be efficiently generated both analytically and numerically. Then applying the same ideas as in Bajnok et al. (2026b), I show that the exponentially small corrections can be easily generated from the perturbative functions with the simple rule:
@@
The advantage of writing the transseries as (52) instead of (9), is that the exponential scales are in one-to-one correspondence with the zeros of the relevant symbol (11). Moreover, the exponentially small corrections are at most first order in every scale $e^{-8\pi gx_{l}^{+}}$ and $e^{-8\pi gx_{j}^{-}}$, and the corresponding $1/g$ expansions can be easily obtained from the perturbative sector with the simple rule in (16). This is the same structure as was studied in Bajnok et al. (2026b). Therefore, (52) puts the observables in points 2. and 3. defined through the matrix Bessel kernel with different values of the mixing angle $\alpha$ in the universal strong coupling structure of Bajnok et al. (2026b). It also provides a more natural framework for describing the resurgence properties of the above-mentioned observables at large values of the ’t Hooft coupling.
According to the results of this paper together with Bajnok et al. (2026b), there is a simple underlying structure in the strong coupling expansion for certain observables of $\mathcal{N}=4$ SYM, at least for those given by the determinant (1). It is an open question whether the same structure holds for other observables in $\mathcal{N}=4$ SYM as well, or physical quantities in different but related models, such as the $O(6)$ model Polyakov (1980); Korchemsky and Radyushkin (1987); Alday and Maldacena (2007). It would also be interesting to relate the obtained transseries structure to the results of Bajnok et al. (2023, 2025a) where a complete resurgence analysis was carried out for a family of generalized energy densities in the $O(N)$ models.
@@ 2 Strong coupling expansion
$\mathcal{I}_{n}(a)=(-1)^{n}\mathcal{I}_{n}(-a)$ (21)
The moments $\mathcal{I}_{n}$ depend on the mixing angle $a$. However, for reasons explained later, I suppress their $a$ dependence and indicate it explicitly only when necessary.
@@
$\mathcal{Z}^{(0,1)}(g|a)=\mathcal{Z}^{(1,0)}(g|-a)\,$, (28)
where the change $a\to-a$ applies to the full dependence on $a$. The same property for the corresponding exponential scales $\Lambda^{2}_{-}$ and $\Lambda^{2}_{+}$. Therefore the prefactor $P^{(0,1)}$ and the coefficients $f^{(0,1)}_{k}$ can be easily obtained from $P^{(1,0)}$ and $f^{(1,0)}_{k}$ by changing $a\to-a$.
@@
$\mathcal{Z}^{(n,m)}(g|a)=\mathcal{Z}^{(m,n)}(g|-a)\,.$ (34)
where the change $a\to-a$ applies again to the full dependence on $a$. The exponential scales in (10) also respect this symmetry.
@@
$\mathcal{I}_{n}(a)=(-1)^{n}\mathcal{I}_{n}(-a)=\mathcal{I}_{n}(a+1)=(-1)^{n}\mathcal{I}_{n}(1-a)\,.$ (40)
@@
$\mathcal{I}_{n}(a+m)\equiv\mathcal{I}_{n}(a)\,$, (41)
@@
That is, the large-$g$ expansion is the same as in the principal branch with the same moments given in (20), but the explicit dependence on $a$ is shifted according to the branch of $a$. To distinguish it from the full $a$ dependence (such as in (34)), I will use the above notation to indicate that the shift acts only on the explicit $a$ dependence and the moments treated as independent quantities. Later we will see, that the same property appears in the relation connecting the perturbative series with the non-perturbative corrections.
One final remark is that although it is formally possible to determine the transseries (9) up to arbitrary order in $1/g$ and $\Lambda_{\pm}^{2}$ with our previous method established in Bajnok et al. (2025b), however, due to the scaling (36), beyond a certain exponential order, it becomes difficult to gain analytical results for the strong coupling coefficients. Following the idea of Bajnok et al. (2026b), in the next section I will show, that (9) can be recast into a new form, in which a natural connection between the perturbative and non-perturbative contributions emerges, making the computation of non-perturbative corrections extremely efficient.
@@ 3 The transseries structure from a different angle
The reason for the cancelation of terms that do not satisfy the condition in (35) remained unknown, as it does not follow trivially from the method presented in Bajnok et al. (2025b). It can also be seen that, by applying certain transformations to the parameter $a$ and the integrals $\mathcal{I}_{n}$, the non-perturbative functions $\mathcal{Z}^{(n,m)}(g)$ with $(n,m)\neq(0,0)$ can easily be obtained from the perturbative part $\mathcal{Z}^{(0,0)}(g)$. In the following I will show that the procedure of Bajnok et al. (2026b) can be generalized to the matrix Bessel kernel (4) with symbol (7), and these properties follow automatically from the analytic structure of the modified symbol (11) and its successive redefinition at each exponential level.
@@
From the method of Bajnok et al. (2025b) and the structure studied in Bajnok et al. (2026b), we expect the appearance of non-perturbative contributions at orders of $e^{-8\pi gx^{\pm}_{j}}$. If we look at the exponential scales $\Lambda_{-}^{2n}\Lambda_{+}^{2m}$ multiplying $\mathcal{Z}^{(n,m)}(g)$ in the transseries (9), we find that the exponential prefactors are:
@@
For different values of $n$ and $m$, the exponent contains both +a+a and a-a terms and can be written as finite linear combinations of the zeros $x_{j}^{\pm}$, with each $x_{j}^{\pm}$ appearing with a multiplicity of one. Therefore, together with the findings of Bajnok et al. (2026b), this suggests that in the case of matrix Bessel kernel, non-perturbative contributions contain exponential factors that are at most first order in each parameter $e^{-8\pi gx_{j}^{+}}$ and $e^{-8\pi gx_{j}^{-}}$. Therefore, the exponential orders can be parametrized as:
@@
$S^{(\delta^{+},\delta^{-})}(a)$ $=S^{(\delta^{-},\delta^{+})}(-a)\,$,
$\mathcal{D}^{(\delta^{+},\delta^{-})}(g|a)$ $=\mathcal{D}^{(\delta^{-},\delta^{+})}(g|-a)\,.$ (54)
@@ 3.1 $1/g$ expansion
Now, this function has zeros at $x=2i\pi x_{j}^{-}$ with $j\geq 1$ and poles at $x=2i\pi x^{-}_{0}$ and $x=2i\pi y_{j}$ with $j\geq 0$. These transformation rules are analogous to what we have found in Bajnok et al. (2026b).
@@ 3.3 A practical example
[…]
@@
The perturbative part in (3.3) reproduces the expansion presented in Basso et al. (2008), while the non-perturbative corrections coincide with the results of Basso and Korchemsky (2009) and Dorigoni and Hatsuda (2015). With the method discussed above, any higher order non-perturbative contribution to the cusp anomalous dimension can be easily obtained in an extremely efficient way. The complete strong coupling structure of $\Gamma$ summarized in Bajnok et al. (2026a).
@@ 4.1 Stokes automorphism
$\mathcal{S}_{\pm}\left[\mathcal{D}^{(\{\},\{\})}\right](\tilde{g})=\tilde{g}\int_{0}^{\infty e^{\pm i\epsilon}}dse^{-\tilde{g}s}\mathcal{B}\left[\mathcal{D}^{(\{\},\{\})}\right](s)\,.$ (107)
@@ 4.2 Asymptotic analysis
Notice that the leading asymptotic behavior of $d^{(\{\},\{\})}$, $d^{(\{0\},\{\})}$ and $d^{(\{\},\{0\})}$ with the coefficients given in (4.2) is compatible with results found in Bajnok et al. (2025b) for the functions $\mathcal{Z}^{(0,0)}(g)$, $\mathcal{Z}^{(1,0)}(g)$ and $\mathcal{Z}^{(0,1)}(g)$, however there are sign differences in the coefficients $A_{0}^{+}$ and $A_{0}^{-}$. As a verification of (4.2), I also performed a Richardson extrapolation on the coefficients of the perturbative function. Its asymptotic behavior agreed with (4.2) with the same coefficient $A_{0}^{+}$ as in (4.2). Furthermore, with the method discussed in Appendix G of Bajnok et al. (2026b), I was also able to separate the first subleading cut of $\mathcal{B}\left[\mathcal{D}^{(\{\},\{\})}\right](s)$. I found that around this cut, its asymptotic behavior is given by: